y=(x-c1)2+c2x3==

1
Expert's answer
2020-10-27T19:07:17-0400

We know that "(a \pm b)^2 = a^2 \pm 2ab +b^2". Then

"y = (x -c_1)^2 +c_2 x^3 = c_1^2 - 2c_1 x +x^2 +c_2x^3"